OpenStudy (anonymous):

Please Help! :) Kathi and Robert had a pottery stand at the craft fair. They sold some of the pottery at the original price of $12.50. Then they decreased the price to $8.50. They sold 88 pieces of pottery. How many cost $12.50, and how many cost $8.50?

OpenStudy (anonymous):

@Wriddler

OpenStudy (anonymous):

Theres not enough information to determine, is there any other info?

OpenStudy (anonymous):

Oops! I forgot to say that they made $852

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

and 88 is total?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, so let them sell x pieces at $12.50 and y at $8.50

OpenStudy (anonymous):

x+y=88

OpenStudy (anonymous):

12.5x + 8.5y = 852

OpenStudy (anonymous):

okay so i have to use the substitution method

OpenStudy (anonymous):

12.5x + 8.5y = 852 x + y = 88

OpenStudy (anonymous):

12.5(88-y) + 8.5y = 852

OpenStudy (anonymous):

combine like terms.......

OpenStudy (anonymous):

1100 - 12.5y + 8.5y +852

OpenStudy (anonymous):

combinde again... (let me kno if I make a mistake plz)

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

LEt me just look through what youve done so far

OpenStudy (anonymous):

1100-4y=852

OpenStudy (anonymous):

-4y+852=1100

OpenStudy (anonymous):

-4y=-248

OpenStudy (anonymous):

y=62

OpenStudy (anonymous):

Good :) now you need to get x

OpenStudy (anonymous):

Ok I figured it out! thanks! its 26

OpenStudy (anonymous):

Yep, you need both for your answer I believe but good job, you understand now then?